• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 02:07:21 on December20,2003

    Entry number 114655

    keyword=swing shift summary

    started shift in controlled access.

    Ramped right Q2 and Q3; right Q3 tripped and required a controlled access to reset power supply. Happened again, but third time was the charm. Right arm is cycled and ready for beam.

    Left arm software was flaky. After changing to septum at 6 degrees in top left of Tools screen, left HRS put Q1 to 3500 A and Q2/A3 to 0.00 regardless of the momentum requested. Dipole NMR would not lock.

    See HALOG entry on magnet woes for full story; short story is Pamela K. and Javier were able to get software working but found two hardware problems. Q1 has stuck bit which forces you to skip every other 6 Amps, and dipole NRM compensating coil voltage will not come on meaning it is not possible to get an NMR lock.

    At end of shift had put 50 A into both septa and observed no beam steering.